integration of (e/x)x+(x/e)x

Grade:12th pass

1 Answers

jagdish singh singh
8 years ago
$May be our Question is like that $\bf{\int \left[\left(\frac{x}{e}\right)^x+\left(\frac{e}{x}\right)^x\right]\cdot \ln(x)dx}$